A builder has 8 lots available for sale.
* 6 lots are greater than one acre
* 2 lots are less than one acre
What is the probability that the next three lots sold will be greater than one acre?
a. 3/4 b. 15/28 c. 27/64 d. 5/14

Answer by stanbon(75887)

A builder has 8 lots available for sale.
* 6 lots are greater than one acre
* 2 lots are less than one acre
What is the probability that the next three lots sold will be greater than one acre?
-----
Binomial Problem with n = 3 and p(> 1 acre) = 6/8 = 3/4
-----------------------------------
P(x = 3) = (3/4)^3 = 27/64
===============================
Cheers,
Stan H.
======================
a. 3/4 b. 15/28 c. 27/64 d. 5/14